# On the geometry of spaces of filtrations on local rings

### v1

@@ 2.4. Log canonical thresholds and normalized volumes
We will write $\mathrm{lct}(\mathfrak{a}_{\bullet})$ when there is no ambiguity. For more properties of lct, we refer to [JM12]. We refer to [LLX20, Zhu23] for a comprehensive survey on normalized volumes.
@@
(c.f. [Zhu24, Lemma 3.4]) There exists $c_{0}=c_{0}(n)>0$ depending only one $n$ such that for any $n$-dimensional klt singularity $x\in X$ and $v\in\mathrm{Val}_{X,x}$,
$\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})$ (2.8)
@@ 2.6. Metric spaces
A point $z$ in a metric space $(S,d)$ is said to be between two points $x$ and $y$, if all three points are distinct, and $d(x,y)=d(x,z)+d(z,x)$.
@@ 2.7. Space of valuations and filtrations
  • the space of valuations with finite log discrepancy, when $x\in X=\mathrm{Spec}R$ is a klt singularity over a field of characteristic 00, denoted by $\mathrm{Val}^{<+\infty}_{X,x}$, and
@@
$\mathrm{DivVal}_{X,x}\subset\mathrm{QMVal}_{X,x}\subset\mathrm{Val}^{<+\infty}_{X,x}\subset\mathrm{Val}^{+}_{X,x}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{h}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ (2.9)
@@ 2.7.1. Weak topologies on $\mathrm{Fil}$
There are more ways to define a topology on the space $\mathrm{Fil}$ (and $\mathrm{Fil}^{s}$ and $\mathrm{Fil}^{s}$, respectively), depending on the point of view.
For example, a filtration $\mathfrak{a}_{\bullet}$ acts naturally on the ring $R$ as a function $\chi_{\mathfrak{a}_{\bullet}}$ (or dually). The product topology, or the “weak-* topology” induced by this action will be called the weak topology, which is the coarest topology such that for any $f\in R\backslash\{0\}$, the function $\phi_{f}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ is continuous. In other words, a subbase of the weak topology is given by $\{\phi_{f}^{-1}(U)\mid U\subset\mathbb{R}_{\geq 0}\text{ is open, }f\in R\backslash\{0\}\}$. Under this topology, all valuation spaces in (2.9) equipped with the weak topology, defined in [JM12], are subspaces of $\mathrm{Fil}$. It is not hard to check that a sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ if and only of $\chi_{k}(f)\to\chi(f)$ for any $f\in R\backslash\{0\}$.
Similarly, a valuation has a natural action on the spaces of filtrations (or dually), that is, any $v\in\mathrm{Val}_{X,x}$ gives a function $\mathrm{ev}_{v}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto v(\mathfrak{a}_{\bullet})$. Hence one can consider the “weak topology” induced by this action, that is, the coarest topology on $\mathrm{Fil}^{s}$ such that $\mathrm{ev}_{v}$ is continuous for any $v\in X^{{}^{\mathrm{an}}}$, $\mathrm{Val}_{X,x}$, $\mathrm{Val}^{+}_{X,x}$ or $\mathrm{DivVal}_{X,x}$, respectively. Similarly one can check that a sequence $\{\mathfrak{a}_{\bullet,k}\}$ converges to $\mathfrak{a}_{\bullet}$ under these topologies if any only if $v(\mathfrak{a}_{\bullet,k})\to v(\mathfrak{a}_{\bullet})$ for any $v$ in the corresponding valuation space. In this paper, we will mainly consider the topology defined using $\mathrm{Val}^{+}_{X,x}$, and call it the $+$-topology. The product topology defined using $\mathrm{DivVal}_{X,x}$ as above was considered in [BdFFU15] for $b$-divisors. We will thus call it the coefficientwise topology.
@@
Since $v(\mathfrak{a}_{\bullet})=v(\widetilde{\mathfrak{a}}_{\bullet})$ by [BLQ24, Proposition 3.9] for $v\in\mathrm{DivVal}_{X,x}$, the functions $\mathrm{ev}_{v}$ do not distinguish $\mathfrak{a}_{\bullet}$ and its saturation. Hence the $+$-topology (and the coefficientwise topology) only makes sense on $\mathrm{Fil}^{s}$. Similar issues appear if we define the topology on $\mathrm{Fil}$ using other valuation spaces. But all the topologies are well-defined on $\mathrm{Fil}^{s}$.
@@ 3.1. Darvas metrics and multiplicities
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}):=\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}).$ (3.1)
@@
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}):=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{b}_{\bullet}).$ (3.2)
@@
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.3)
@@
$\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})\leq\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})$ (3.4)
@@
Using the identification between $\mathrm{Fil}$ and $\mathcal{N}$ Lemma 2.8, we get a pseudometric $d_{1}$ on $\mathcal{N}$. The following lemma asserts that $(\mathrm{Fil},d_{1})$ is a rooftop metric space in the sense of [Xia23], which is an analogue of [Dar19, Proposition 3.35] and [BJ21, Lemma 3.2].
@@
$d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$ (3.5)
@@ 3.2. Measures and geodesics
The goal of this section is to justify the term geodesic, in the sense that it computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. In other words, it is a distance-minimizing continuous curve with respect to the metric $d_{1}$.
We first give a formula for the geodesic between two saturated filtrations, and use it to show that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ when $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$..
@@
$\mathfrak{a}_{\bullet,t}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{c_{t,v}}),$ (3.6)
@@ 3.2.1. Distance-minimizing property
$\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t}.$ (3.7)
@@
$d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}}).$ (3.8)
@@
$\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda}.$ (3.9)
@@ 3.2.2. Duistermaat-Heckman measures and continuity
In this section $R$ contains a field $\mathbbm{k}$. We briefly recall the measure $\mu$ introduced in [BLQ24, Section 4] encoding multiplicities and its properties.
@@
The first two statements are [BLQ24, Proposition 4.7] and the last two are [BLQ24, Proposition 4.9]
@@
Fix $(a,b)\in\mathbb{R}^{2}$. Let $A:=[a+\alpha]\times[b+\beta]\subset\mathbb{R}^{2}$ be a rectangle, where $\alpha,\beta\in\mathbb{R}_{>0}$. Then there exists $M:=M(a,b)\in\mathbb{R}_{>0}$ such that $\mu(A)\leq M\alpha\beta$ when $\alpha,\beta\leq 1$.
@@
where the second and the third equality follow from Lemma 3.8(3), and the first inequality follows from Lemma 3.8(2). Hence proof is finished with $M:=n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$. ∎
@@
Suppose $R$ contains a field. Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the map $[0,1]\to\mathrm{Fil}$, $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous, where $[0,1]$ is equipped with the Euclidean topology and $\mathrm{Fil}$ is equipped with the topology induced by the pseudometric $d_{1}$.
@@
$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})=\mu(\Delta_{1}\backslash\Delta_{2})+\mu(\Delta_{2}\backslash\Delta_{1}).$ (3.10)
First assume that $10<t_{1}<t_{2}<$ 1. By Lemma 3.8(4) we have
@@
Note that $(x,y)\in\Delta$ satisfies $x\leq 1$ and $y\leq D$. So $\Delta$ can be covered by finitely many rectangles $R_{j}$ of the form $[a_{j},a_{j}^{\prime}]\times[b_{j},b_{j}^{\prime}]$, satisfying
@@
$d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1}).$ (3.11)
@@
$d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n}(2^{n-1}(D-1)^{2}+1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}.$ (3.12)
@@ 3.3. Continuity along monotonic sequences
In this section we prove some convergence result along monotonic results.

### v2

@@ 2.4. Log canonical thresholds and normalized volumes
We will write $\mathrm{lct}(\mathfrak{a}_{\bullet})$ when there is no ambiguity.
For more properties of lct, we refer to [JM12]. We refer to [LLX20, Zhu25] for a comprehensive survey on normalized volumes.
@@
(c.f. [Zhu24, Lemma 3.4]) There exists $c_{0}=c_{0}(n)>0$ depending only on $n$ such that for any $n$-dimensional klt singularity $x\in X$ and $v\in\mathrm{Val}_{X,x}$,
(2.8) $\frac{A_{X}(v)}{v(\mathfrak{a}_{\bullet})}\leq c_{0}\cdot\frac{\widehat{\mathrm{vol}}_{x,X}(v)}{\widehat{\mathrm{vol}}(x,X)}\cdot\mathrm{lct}(X;\mathfrak{a}_{\bullet})$
@@ 2.6. Metric spaces
A point $z$ in a metric space $(S,d)$ is said to be between two points $x$ and $y$, if all three points are distinct, and $d(x,y)=d(x,z)+d(z,y)$.
@@ 2.7. Space of valuations and filtrations
  • the space of valuations with finite log discrepancy, when $x\in X=\mathrm{Spec}R$ is a klt singularity over a field of characteristic 0, denoted by $\mathrm{Val}^{<+\infty}_{X,x}$, and
@@
(2.9) $\mathrm{DivVal}_{X,x}\subset\mathrm{QMVal}_{X,x}\subset\mathrm{Val}^{<+\infty}_{X,x}\subset\mathrm{Val}^{+}_{X,x}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{h}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$
@@ 2.7.1. Weak topologies on $\mathrm{Fil}$
There are more ways to define a topology on the space $\mathrm{Fil}$ (as well as on $\mathrm{Fil}^{s}$ and $\mathrm{Fil}^{h}$), depending on the point of view.
For example, a filtration $\mathfrak{a}_{\bullet}$ acts naturally on the ring $R$ as a function $\chi_{\mathfrak{a}_{\bullet}}$ (or dually). The product topology, or the “weak-* topology” induced by this action will be called the weak topology, which is the coarsest topology such that for any $f\in R\backslash\{0\}$, the function $\phi_{f}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)$ is continuous. In other words, a subbase of the weak topology is given by $\phi_{f}^{-1}(U)\mid U\subset\mathbb{R}_{\geq 0}\text{ is open, }f\in R\backslash\{0\}\}$. Under this topology, all valuation spaces in (2.9) equipped with the weak topology, defined in [JM12], are subspaces of $\mathrm{Fil}$. It is not hard to check that a sequence $\{\mathfrak{a}_{\bullet,k}\}\subset\mathrm{Fil}$ converges weakly to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}$ if and only if $\chi_{k}(f)\to\chi(f)$ for any $f\in R\backslash\{0\}$.
Similarly, a valuation has a natural action on the spaces of filtrations (or dually), that is, any $v\in\mathrm{Val}_{X,x}$ gives a function $\mathrm{ev}_{v}:\mathrm{Fil}\to\mathbb{R}_{\geq 0}$, $\mathfrak{a}_{\bullet}\mapsto v(\mathfrak{a}_{\bullet})$. Hence one can consider the “weak topology” induced by this action, that is, the coarsest topology on $\mathrm{Fil}^{s}$ such that $\mathrm{ev}_{v}$ is continuous for any $v\in X^{{}^{\mathrm{an}}}$, $\mathrm{Val}_{X,x}$, $\mathrm{Val}^{+}_{X,x}$ or $\mathrm{DivVal}_{X,x}$, respectively. Similarly, one can check that a sequence $\{\mathfrak{a}_{\bullet,k}\}$ converges to $\mathfrak{a}_{\bullet}$ under these topologies if and only if $v(\mathfrak{a}_{\bullet,k})\to v(\mathfrak{a}_{\bullet})$ for any $v$ in the corresponding valuation space. In this paper, we will mainly consider the topology defined using $\mathrm{Val}^{+}_{X,x}$, and call it the ++-topology. The product topology defined using $\mathrm{DivVal}_{X,x}$ as above was considered in [BdFFU15] for $b$-divisors. We will thus call it the coefficientwise topology.
@@
[…]
@@ 3.1. Darvas metrics and multiplicities
(3.1) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})=d_{1}(\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}):=\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}).$
@@
(3.2) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet}):=d_{1}(\mathfrak{a}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{b}_{\bullet}).$
@@
(3.3) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})+d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$
@@
(3.4) $\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{c}_{\lambda})\leq\ell(R/\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})+\ell(R/\mathfrak{b}_{\lambda}\cap\mathfrak{c}_{\lambda})-\ell(R/\mathfrak{b}_{\lambda})$
@@
Using the identification between $\mathrm{Fil}$ and $\mathcal{N}$ Lemma 2.8, we get a pseudometric $d_{1}$ on $\mathcal{N}$. The following lemma asserts that $(\mathrm{Fil},d_{1})$ is a rooftop metric space in the sense of [Xia23], which is an analogue of [Dar19, Proposition 3.35] and [BJ25, Lemma 3.2].
@@
(3.5) $d_{1}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet},\mathfrak{a}_{\bullet}\cap\mathfrak{c}_{\bullet})\leq d_{1}(\mathfrak{b}_{\bullet},\mathfrak{c}_{\bullet}).$
@@ 3.2. Measures and geodesics
The goal of this subsection is to justify the term geodesic, in the sense that it computes the distance between $\mathfrak{a}_{\bullet,0}$ and $\mathfrak{a}_{\bullet,1}$. In other words, it is a distance-minimizing, continuous curve with respect to the metric $d_{1}$.
We first give a formula for the geodesic between two saturated filtrations, and use it to show that $\mathfrak{a}_{\bullet,t}\in\mathrm{Fil}^{s}$ when $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}^{s}$.
@@
(3.6) $\mathfrak{a}_{\bullet,t}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{c_{t,v}}),$
@@ 3.2.1. Distance-minimizing property
(3.7) $\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}\subset\mathfrak{a}_{\lambda,t}.$
@@
(3.8) $d_{1}(a_{\bullet,0},\mathfrak{a}_{\bullet,t^{\prime}})=d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t})+d_{1}(\mathfrak{a}_{\bullet,t},\mathfrak{a}_{\bullet,t^{\prime}}).$
@@
(3.9) $\mathfrak{a}_{\lambda,t}/\mathfrak{b}_{\lambda}\cong\mathfrak{c}_{\lambda}/\mathfrak{b}^{\prime}_{\lambda}.$
@@ 3.2.2. Duistermaat-Heckman measures and continuity
In this subsection, we assume that $R$ contains a field $\mathbbm{k}$. We first briefly recall the measure $\mu$ introduced in [BLQ24, Section 4] encoding multiplicities and its properties.
@@
The first two statements are [BLQ24, Proposition 4.7] and the last two are [BLQ24, Proposition 4.9]. ∎
For a rectangle, the measure $\mu$ is controlled by the Lebesgue measure.
@@
Fix $(a,b)\in\mathbb{R}^{2}$. Let $A:=[a,a+\alpha]\times[b,b+\beta]\subset\mathbb{R}^{2}$ be a rectangle, where $\alpha,\beta\in\mathbb{R}_{>0}$. Then there exists $M:=M(a,b)\in\mathbb{R}_{>0}$ such that $\mu(A)\leq M\alpha\beta$ when $\alpha,\beta\leq 1$.
@@
where the second and the third equality follow from Lemma 3.8(3), and the first inequality follows from Lemma 3.8(2). Hence the proof is finished with $M:=n^{2}(a+1)^{n-1}(b+1)^{n-1}\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})$. ∎
@@
Suppose $R$ contains a field. Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}$, the map $0,1]\to\mathrm{Fil}$, $t\mapsto\mathfrak{a}_{\bullet,t}$ is Lipschitz continuous, where $[0,1]$ is equipped with the Euclidean topology and $\mathrm{Fil}$ is equipped with the topology induced by the pseudometric $d_{1}$.
@@
(3.10) $d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})=\mu(\Delta_{1}\backslash\Delta_{2})+\mu(\Delta_{2}\backslash\Delta_{1}).$
First assume that $0<t_{1}<t_{2}<1$. By Lemma 3.8(4) we have
@@
Note that $(x,y)\in\Delta$ satisfies $x\leq 1$ and $y\leq D$. So $\Delta$ can be covered by finitely many rectangles $R_{j}$ of the form $a_{j},a_{j}^{\prime}]\times[b_{j},b_{j}^{\prime}]$, satisfying
@@
(3.11) $d_{1}(\mathfrak{a}_{\bullet,t_{1}},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n+1}(D+1)^{n-1}D(D-1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})(t_{2}-t_{1}).$
@@
(3.12) $d_{1}(\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,t_{2}})\leq n^{2}2^{n}(2^{n-1}(D-1)^{2}+1)\mathrm{e}(\mathfrak{a}_{\bullet,0}\cap\mathfrak{a}_{\bullet,1})t_{2}.$
@@ 3.3. Continuity along monotonic sequences
In this subsection we prove some convergence results along monotonic sequences. The results are partially inspired by the study of the completion of certain metric spaces in the global setting, for example, [Dar17, Xia23, BJ25].